#9. The normal human body has an average core or interior temperature of 37C. At a skin temperature of 34 C, how much heat is removed through conduction? Assume a thermal conductivity of unperfused (i.e. bloodless) tissue of 18 (Cal x cm)/(m^2 x hr x C), 3 cm thickness of the tissue between the interior and the exterior of the body and 1.5 m2 total surface area of the human body. 

Choices: 

  1. 36 Cal/hr
  2. 29 Cal/hr
  3. 27 Cal/hr
  4. 18 Cal/hr
  5. 22.5 Cal/hr


#10. If the human body needs to get rid of 150 Cal of heat per hour through conduction alone, at what temperature should the skin be if a core temperature of 37C is maintained? Assume a thermal conductivity of unperfused (i.e. bloodless) tissue of 18 (Cal x cm)/(m^2 x hr x C), 3 cm thickness of the tissue between the interior and the exterior of the body and 1.5 m2 total surface area of the human body. 

  1. 20.333 C  
  2. 20.889 C 
  3. 22 C 
  4. 19.778 C 
1
Expert's answer
2021-01-01T14:06:54-0500

#9 "q=\lambda\cdot\frac{\Delta T}{\Delta x}\cdot A=18\cdot\frac{37-34}{3}\cdot 1.5=27(Cal/hr)" . Answer


#10 "q=\lambda\cdot\frac{\Delta T}{\Delta x}\cdot A\to t=37-\frac{q\cdot\Delta x}{\lambda\cdot A}="


"=37-\frac{150\cdot3}{18\cdot 1.5}=20.333(°C)" . Answer